OpenStudy (anonymous):

Help!! I'm willing to give medals! Vinny is serving punch to guests at a party. He has \frac{2}{3} of a gallon of punch, and he pours \frac{3}{64} of a gallon into each guest's cup, filling each to the top. At the most, how many guests' cups can he fill to the top?

OpenStudy (perl):



OpenStudy (anonymous):

\[\frac{ 3 }{ 64}\times x =2/3\]

OpenStudy (anonymous):

ok